OpenStudy (anonymous):

how would i solve this ? anyone know the steps ? 15^2 - 15t -30 = 0

OpenStudy (anonymous):

factor ?

OpenStudy (skullpatrol):

Factor out 15 first

OpenStudy (unofficialllyy):

Exponents first. Use PEMDAS

OpenStudy (unofficialllyy):

Parenthesis, Exponents, Multiplication, Division, Addition, then Subtraction.

geerky42 (geerky42):

@kisntell16 You mean \(15t^2 - 15t -30 = 0\), right?

OpenStudy (anonymous):

so i got 3t^ - 5t- 5=0 from factoring 15

OpenStudy (anonymous):

and yes c: @geerky42

OpenStudy (unofficialllyy):

Oh, well then never mind about my answer lol.

OpenStudy (unofficialllyy):

Good luck!

OpenStudy (anonymous):

oh ok but thanks tho c: @Unofficialllyy

OpenStudy (anonymous):

well wait how did you do it tho ? show me maybe im wrong @Unofficialllyy

OpenStudy (skullpatrol):

Factor out 15

OpenStudy (anonymous):

yeah an i got 3t^ -5t-5=0 @skullpatrol

OpenStudy (skullpatrol):

Try again please

OpenStudy (unofficialllyy):

I thought your question was \[15^{2}-15t-30=0\] and if that was the case, you would do 15^2 first, since there are no parenthesis and that is the exponent. But you said the question was different.

OpenStudy (anonymous):

well i divided a 5 out of all of it 5 \15^, and 15t and 5/-30 @skullpatrol

OpenStudy (anonymous):

and no youre right thats the question i was asking ? sorry people were confusing me @Unofficialllyy

OpenStudy (skullpatrol):

Divide each term by 15

OpenStudy (unofficialllyy):

Well first you would do 15^2

OpenStudy (anonymous):

ok well in tht case i got t^-1t-2=0 from that @skullpatrol

OpenStudy (unofficialllyy):

I'm confused. What is the original question?

OpenStudy (anonymous):

the one you put down was the original @Unofficialllyy

OpenStudy (skullpatrol):

Now factor $$\Huge t^2-t-2$$

OpenStudy (anonymous):

it just told me to solve @Unofficialllyy

OpenStudy (anonymous):

how ? @skullpatrol

OpenStudy (unofficialllyy):

I have to go, I'm sorry. Hopefully @skullpatrol can help!

OpenStudy (anonymous):

i hope so too haha ! but thanks for trying c: @Unofficialllyy

OpenStudy (skullpatrol):

Look for factors of -2 that add up to -1.

OpenStudy (anonymous):

i factor and i got (x-2) and (x + 1)

OpenStudy (anonymous):

@skullpatrol

OpenStudy (skullpatrol):

Very good :) Now your equation looks like (x-2)(x+1)=0

OpenStudy (anonymous):

so i did it c:? !!

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

one last step lol

OpenStudy (skullpatrol):

Either x-2=0 or x+1=0

OpenStudy (anonymous):

what do you mean ?? haha @jrain35

OpenStudy (anonymous):

thanyouc: @skullpatrol

OpenStudy (skullpatrol):

The values of x that will make (x-2)(x+1)=0 a true statement. Are 2 and -1.